# CCoE: A Compact LLM with Collaboration of Experts

### v1

@@ Abstract
In the domain of Large Language Model (LLM), LLMs demonstrate significant capabilities in natural language understanding and generation. With the growing needs of applying LLMs on various domains, it is a research question that how to efficiently train and build a model that has expertise in different domains but with a low training cost. We propose CCoE architecture, a framework of easily coupling multiple strong domain experts together to fuse into a big LLM, provides a collective way of utilizing the different domain expert LLMs. Besides, training a large collaborative of multiple expert LLMs requires a high requirements on training sources. CCoE bypasses this problem through isolating other experts and train each expert separately. The design of CCoE assembles multiple expert LLMs through the CoE (Collaboration of Experts) layer. Each CoE layer could have one or more expert LLMs. Expert LLMs have different number of layers and have been well-trained for different domain tasks. Each expert is fine-tuned to be able to achieve the comparable results with SOTA domain LLMs. We start from 5 experts in the domain of Code, Math, Law, text-to-SQL and Medical. The results indicate that our CCoE framework can easily and efficiently boost nearly 10%-20% performance on original base model in different domains but using less resources on training, as well as inference.
@@ 4.1 Design of CCoE
We can denote the CCoE framework in Figure 1 as $E_{b,L},E_{1,l_{1}},E_{2,l_{2}},\cdots,E_{3,l_{3}}$, Assume $E_{b,L}$ is the base LLM with total number of $L$ number of layers. One of expert processes an input text to generates its continuation. Formally, for given an input prompt $x$, the output of CCoE can be expressed as following:
@@ 4.2 Rule-based Gating
4.3 Expert Training
@@ 4.4 Expert LLM Architecture
$argmin_{\mathcal{L}_{f_{\theta}}}=argmin_{\ell_{({\mathcal{L}_{i,t_{k}},\mathcal{L}_{j,t_{k}}}})}\forall i,j\in E,\forall t_{k}\in T\text{ }\textrm{s.t.}\text{ }min\mathcal{L}_{val}$ (8)
@@ 5.1 Training Data
We collect 5 different domain datasets: Math, Code, Finance, law, Medical, and Text-to-SQL. Each domain focuses on enhancing model capabilities by injecting domain knowledge or providing reasoning training to experts. In our training corpus, we mainly used the publicly available and high quality datasets. The data composition is illustrated at Figure 3.
Math. Mathematical reasoning mainly relies on model’s inference capabilities and already lead the various approaches to improve the performance on math-related tasks. In our approach, we selected the recent math training dataset that has been shown to effectively boost model’s math performance. [3]. Besides, using the Chain-of-thought and RLHF technologies can further improve model’s math performance [26] [4]. We also constructed our training dataset by providing a step-by-step solution way rather than the vanilla format, and apply the dpo training to align with the preferences. In the our math dataset, we collected nearly 400K instructions mainly from MetaMathQA [^1] https://huggingface.co/datasets/meta-math/MetaMathQA.
Code. The code generation usually combines the execution results with structured code. These pairs are verified and filtered by code interpreter before being used to fine-tune large language models. To optimize the model’s reasoning capabilities, the data is usually formatted with COT prompting. We adopt the commonly used datasets from code instruction 128K alpaca  [^2] https://huggingface.co/datasets/iamtarun/code_instructions_120k_alpaca, code feedback [27] and MBPP [28]. Our merged dataset includes various programming languages and covers major programming topics. We organized the fine-tuning data into instruction-based format with over 200K number of pairs.
Text-to-SQL. The Text-to-SQL (NL2SQL) is a well-known challenge in NLP and database community. It targets to translate the natural language input into corresponding and valid Structured Query Language (SQL). It requires the LLM to understand the semantic purpose of natural language input, the target database schema and the grammar of SQL. Current proposed solution of this task includes SFT, or craftily designed prompt with paradigm of "prompt" learning or "in-context learning". We explore the SFT paradigm to empower LLM Text-to-SQL capability. We collect our dataset with total of 133K instruction-based pair and mainly from WikiSQL [29], Chase [30], and DuSQl [31].
Medical. The challenges of developing LLM-based medical applications are lack of factual inaccuracies, background knowledge for diagnoses, and insufficient grounding in real-world experience. There are many elaborately collected medical datasets that aim to optimize LLM capabilities in medical domain. We adapt the medical data from MedQA [32], MedicalGPT [33], and HuaTuo [34], and create a over 849K English and Chinese QA-paired medical dataset that mixes English and Chinese.
law. We mainly focus on the Chinese mainland law system because there is a possibility that some cases could have different decisions under the Common Law and Civil Law. We collected the instruction-based Civil Law dataset from some influential projects [35] [36]. Our law dataset includes a total of nearly 300K instructions-based pairs that involved with different tasks: decision prediction in a case, legal Q&A, judgment prediction, judicial reading comprehension, legal case classification, etc.
@@ 5.2 Data processing and training
The quality for open-sourced dataset is critical for training expert models. We firstly use Min-Hash and ’text-embedding-ada-002’ embedding model to remove highly similar questions. In downstream fine-tuning community, domain datasets often mix with pre-train dataset to avoid degradation of model’s general capabilities or they train the model with multiple-stages. However, in CCoE framework, each expert LLM fine-tuning only need to focus on learning the domain data. Our dataset preparation only needs to augment the domain data quality. We aim to overcome this through three dimensions: 1. comprehensive 2. complex 3. diverse. Based on our deliberately explorations on domain datasets, the Math dataset is mainly sampled from MetaMath [3]. It uses the bootstrapping approach to rephrase the math question with enough complexity, diversity and comphrensitiviy.
@@
We also conducted the DPO training on Math, Code domains. To create the direct preference pairs, we ask the expert model to generate 5 outputs including the exact answers and reasoning steps with a sampling temperature of 0.8. We adopt the template form [39] to rank the answers and reasoning steps based on the dimensions of completeness, correctness, and readability. Comparing to directly rank the 5 model outputs, we decide to use the pairwise ranking to have a better stable ranking result. We conducted pairwise ranking among them and manually solved the conflicts if there are multiple outputs have the same ranking position or a loop in the ranking positions. We use the highest ranked output as the chosen answer and lowest ranked output as the rejected answer.
@@ 5.3 Model Settings and Evaluation
We use QWen1.5 [40] 7B and 13B parameters as our backbone models. To simplify our training process, we do not expand the vocabulary size and continue to train the word embedding for each domain. We conducted a comprehensive experiments on hyper-parameters $L$, the number of layer in each expert. We explored the model performance gains when we increase the $L$. In our context, we want to limit the layers of each expert to the upper bound $\left\lfloor\frac{L_{base}}{2}\right\rfloor$ because we want to control the overall CCoE model parameters under the constraints of computational resources. Many existing literature have already demonstrate that the deeper layers of LLM often acquire more difficult concepts and can be easily recognized for complex tasks [25] [41]. We start to fine-tune the layers starting from the last 4 layers and add 4 layers every time. We select the number of layers in each expert that gives the highest performance gains. The results of measuring performance gains are listed in Appendix.
@@
Table 1 shows the testing accuracy of our 5 domain experts on GSM8K [^3] https://paperswithcode.com/sota/arithmetic-reasoning-on-gsm8k, MBPP [^4] https://paperswithcode.com/sota/code-generation-on-mbpp, Human Eval [^5] https://paperswithcode.com/sota/code-generation-on-humaneval, LawBench  [^6] https://lawbench.opencompass.org.cn/leaderboard, MedQA [^7] https://paperswithcode.com/sota/question-answering-on-medqa-usmle, and Spider [^8] https://yale-lily.github.io/spider benchmarks. Our CCoE framework is able to achieve the comparable performance and have the greatly outperformed the baseline model Qwen1.5 7B chat and 14B chat [^9] https://qwenlm.github.io/blog/qwen1.5/. Our results illustrate that there are nearly 10%-20% improvement on the baseline models in 5 domains. This indicates that the CCoE framework can easily boost the base model performance in different downstream domains. However, expert in Math has the less improvements after fine-tuning (CCoE 14B Math). We think the base model already achieves a comparably high test results on the benchmark. Our current training strategies can only play limited contributions.
To answer the RQ2, Law and medical domains need to have more fine-tuning layers to delivery better test results. We think this is due to that they have less portion in pre-training corpus and some of their knowledge are contemporary. To revise the contemporary knowledge requires to correct the knowledge neurons. Having more fine-tuning layers allows us to modify the knowledge neurons in the lower layers. However, Math and Code may have larger portions during the pre-training and already have enough knowledge. Our current fine-tuning emphasizes on improving model’s reasoning capability through CoT, question bootstrapping.[3], and self-verification [6].
@@
The Text-to-SQL requires the LLM to understand the SQL syntax, database schema, and the capability of LLM is also very essential. We observed that the Qwen1.5 base model can give the some exact and partial correct results with few-shot prompt. However, they fail to generate many correct SQLs, especially for those hard-level questions. Besides, even though we have conducted the fine-tuning, there is still a performance gap between the models on Spider Leader-board. Comparing to the base model, there is a big improvement. We only conducted limited study on the effect of prompting in Text-to-SQL. In the future, we plan to continue to study the In-Context Learning and calibration of hints to improve the Text-to-SQL performance.
To answer the RQ 3, We select the open-sourced LLaMA2 7B chat model [42] as our comparison model. LLaMA2 7B has the same model scale as the QWen1.5 7B model. To exclude other confounding variables, we train the LLaMA2 7B chat model with same number of expert layers and same dataset. From table 4, we can see that LLaMA model is able to receive a higher performance boost than Qwen1.5 base model. We speculate that a comparatively low performance base model can easily to be enhanced through the downstream fine-tuning and data augmentation strategies. However, LLaMA-based experts still receive low results in 5 domain evaluations. We think this is because expert training requires to utilize the lower layers of the base model to extract the language or even semantic features, which are still decided by the base model’s capabilities. We can conclude that the CCoE framework can boost the performance of base model in different downstream domains regardless of model’s family but expert’s performance is constraint on its base model’s capability. We also realize the scaling law applies to the expert training. It is difficult to continue to improve expert’s performance by increasing the number of training data without changing the expert size.
@@ 6 Summary and Future Work
In this paper, we propose the CCoE framework that aims to boost the base model’s overall performance in all perspectives through utilizing the idea of collaborating multiple of experts (CoE) LLM layers. All of the expert LLM layers are coupled at the base model and becomes the expert sub-network. We conducted the experiments on Math, Code, Law, Medical and Text-to-SQL domains. Our results indicate that there are at nearly 10%-20% improvements on these 5 downstream domains. This proves that CCoE framework is able to boost the overall performance of the base model with the advantages of low-cost, high-interpretability and efficiency. Moreover, CCoE also has the high extensibility through scaling up the number of expert layers to fulfill more open-ended tasks. However, the current routing of the expert layers depends on the API endpoints trace. In the future, We plan to use the LLM to help us route the prompt to the correct expert sub-network. We can extend the prompt’s content to ask the base model to classify the prompt’s task or domain during the first generation and then run the second inference to route the real prompt to the correct expert sub-network. We understand this could result in the high Time to First Token (TTFT), but there is still great potential of the CCoE network for us to explore.

### v3

@@ Abstract
In the domain of Large Language Model (LLM), LLMs demonstrate significant capabilities in natural language understanding and generation. With the growing needs of deploying LLMs across various domains, it is a research question that how to efficiently train and build a model that has expertise in different domains but with a low training cost. We propose the CCoE architecture, a framework that facilitates the seamless integration of multiple specialized domain experts into a unified large LLM, offers a collective approach to leveraging domain experts. Besides, training a large LLM with collaborative of multiple experts requires a high requirements on training sources. CCoE bypasses this problem through isolating other experts and train each expert separately. The design of CCoE assembles multiple experts through the CoE (Collaboration of Experts) layer. Each CoE layer could have one or more experts. Experts in the CCoE are loosely coupled, which allows for scalability to efficiently accommodate any number of experts. Each expert could have different number of layers and is fine-tuned to be able to achieve the comparable results with SOTA domain LLMs. We start from 5 experts in the domains of Code, Math, Law, Text-to-SQL and Medical. The results indicate that our CCoE framework can easily and efficiently boost performance by nearly 10%-20% on the original base model across different domains, while using significantly fewer resources in training, as well as inference.
@@ 4.1 Design of CCoE
We can denote the CCoE framework in Figure 1 as $E_{b,L},E_{1,l_{1}},E_{2,l_{2}},\cdots,E_{n,l_{n}}$, Assume $E_{b,L}$ is the base LLM with total number of $L$ number of layers. One of expert processes an input text to generates its continuation. Formally, for given an input prompt $x$, the output of CCoE can be expressed as following:
@@
CCoE has the advantage of compressing the model to reduce physical model size through the shared layers. For each prompt query, CCoE only needs to activate one expert to generate the answer. This advantage becomes more obvious if we scale up the number of experts. We define the activation ratio as the number of activated parameters divide by the total number of parameters of CCoE model. Take a transformer-based model without bias as an example. Assume the vocabulary size is $V$, hidden size is $d$. The activation ratio can be estimated as following:
$ActivationRatio=\frac{2*Vd+12*L*d^{2}}{2*Vd+12*L*d^{2}+n*Vd+\sum_{i=1}^{n}l_{i}*12d^{2}}$ (8)
As we scale up the number of experts, the numerator is constant and denominator is increasing. The activation ratio is becoming smaller (Figure 2), which represents that the CCoE only needs to activate a small portion of parameters during the inference.
Figure 2: The change of activation ratio with different number of experts in 7B, 14B, 32B, and 72B base models with 4 and 8 expert layers
@@ 4.2 Rule-based Gating
4.3 CCoE Training
In downstream fine-tuning community, domain datasets often mix with pre-train dataset to avoid degradation of model’s general capabilities.
@@ 4.4 Expert LLM Architecture
$argmin_{L_{f_{\theta}}}=argmin_{{L_{i,t_{k}}}}\forall i\in E,\forall t_{k}\in T\text{ }\textrm{s.t.}\text{ }min\mathcal{L}_{val}$ (9)
@@ 5.1 Training Data
We collect 5 different domain datasets: Math, Code, Finance, law, Medical, and Text-to-SQL. In our training corpus, we mainly used the open-sourced and high quality datasets. The data composition is illustrated at Figure 3.
Math. Mathematical reasoning mainly relies on model’s inference capabilities and already lead the various approaches to improve the performance on math-related tasks. In the our math dataset, we collected nearly 400K instructions mainly from MetaMathQA [^1] https://huggingface.co/datasets/meta-math/MetaMathQA.
Code. We adapt the commonly used datasets from code instruction 128K alpaca  [^2] https://huggingface.co/datasets/iamtarun/code_instructions_120k_alpaca, code feedback [26] and MBPP [27]. Our merged dataset includes various programming languages and covers major programming topics. We organized the fine-tuning data into instruction-based format with over 200K number of pairs.
Text-to-SQL. Text-to-SQL (NL2SQL) dataset usually includes the databse tables, databse schema and QA pairs. We formed dataset with total of 133K instruction-based pair and mainly from WikiSQL [28], Chase [29], and DuSQl [30].
Medical. Our medical training dataset is mainly from MedQA [31], MedicalGPT [32], and HuaTuo [33]. It consists over 849K English and Chinese QA-paired medical dataset that mixes English and Chinese.
law. We focus on the instruction-based Civil Law dataset from some influential projects [34] [35]. Our law dataset includes a total of nearly 300K instructions-based pairs that focuses on the Chinese mainland law system.
@@ 5.2 Data processing and training
Expert Training. Many researches have investigated different training strategies to bolster model’s zero-shot performance in different domains. Domain datasets usually consists of diverse tasks, various question formats, and vast amount of knowledge. Common SFT approaches include the multi-task training, sequential training, mixed data training, and so on. However, cross-domain training can easily have the task conflicts that greatly hurt model’s learning in different domains. Sequential training suffers the problem of catastrophic forgetting of the domain knowledge, which learned in the early stage. Mixed data training requires great amount of efforts on studying the composition of training data. In CCoE framework, we utilize the idea of "divide-and-conquer". Instead of letting single LLM to enhance its capabilities in 5 domains, CCoE constructs 5 learning tasks for 5 experts. It creates individual expert that is responsible for learning one domain or task. This is equivalent to train 5 different domain LLMs that only need to be specialized in their own area. There is no interleaved learning tasks among the experts and avoid the mutual interference in multi-task learning. Since CCoE creates a new LLM for each domain, it creates a copy of base model’s parameters and only update on the copy. CCoE does not update any parameters in base model. The training of CCoE experts would not harm the general capabilities of the base model. The expert training paradigm in CCoE directly resolve the challenges and greatly simplifies the procedure of fine-tuning LLM in multiple domains.
The effectiveness of expert training in CCoE is another issue that we need to justify. In real world, the limitation of computing resources prevents us to run multiple specialized domain LLMs together. CCoE framework aims to provide a lightweight approach that assembles multiple domain LLMs into the base LLM. To address the constraints of resources, CCoE proposes to create expert layers through only fine-tuning several layers of base LLM with domain data. The expert layers can form an individual and complete LLM with the shared layers in base LLM. CCoE intuitively selects the upper layers of LLM as the expert layers. We have discussed the KN thesis previously that Knowledge Neurons store the factual relations. However, many existing literature has already demonstrate the upper layers of LLM often possess difficult concepts and more context knowledge that can be easily recognized for complex tasks [25] [39]. We examined our training corpus that contains a diverse format of questions including text summarization, cloze test, entity extraction, sentence correction, etc. All of these requires a strong understanding and reasoning abilities that are unlikely to be resolved through the lower layers of the base LLM. We believe that the CCoE framework provides a feasible and efficient way of transferring one LLM into multiple specialized LLMs.
@@ 5.3 Model Settings and Evaluation
We use QWen1.5 [40] 7B and 13B parameters as our backbone models. We trained and assembled 5 domain experts into two CCoE models on 8 NVIDIA H800 80G GPUs. To simplify our training process, we do not expand the vocabulary size and continue to train the word embedding for each domain. We conducted a comprehensive experiments on hyper-parameters $L$, the number of layer in each expert. We explored the model performance gains when we increase the $L$. In our context, we want to limit the layers of each expert to the upper bound $\left\lfloor\frac{L_{base}}{2}\right\rfloor$ because we want to control the overall CCoE model parameters under the constraints of computational resources. In order to answer RQ1, we start to fine-tune the layers starting from the last 4 layers and add 4 layers every time. We select the number of layers in each expert that gives the highest performance gains. The results of measuring performance gains are listed in Appendix.
Table 1: The metrics of CCoE-7B base and CCoE-13B base with 5 experts. The Training Time ratio takes the amount of time needed to train the base model 7B and 13B in 5 domains as the base: $\frac{Training-time_{CCoE}}{Training-time_{Base-Model}}$. We measured the change of avg. Time-to-First-Token on various prompt length (20-512) sampled from each domain.
. Metric CCoE-7B base CCoE-14B base # total params 20.7B 40.7B # Activated Params 7B 14B Activation Ratio 0.33 0.34 Training Time Ratio 0.76 0.73 $\Delta$ TTFT (Time-to-First-Token) +0.117e-4 0.1842e-4
@@
Table 1 shows the testing accuracy of our 5 domain experts on GSM8K [^3] https://paperswithcode.com/sota/arithmetic-reasoning-on-gsm8k, MBPP [^4] https://paperswithcode.com/sota/code-generation-on-mbpp, Human Eval [^5] https://paperswithcode.com/sota/code-generation-on-humaneval, LawBench  [^6] https://lawbench.opencompass.org.cn/leaderboard, MedQA [^7] https://paperswithcode.com/sota/question-answering-on-medqa-usmle, and Spider [^8] https://yale-lily.github.io/spider benchmarks. Our CCoE framework is able to achieve the comparable performance and have the greatly outperformed the baseline model Qwen1.5 7B chat and 14B chat [^9] https://qwenlm.github.io/blog/qwen1.5/. For each domain, the result illustrates that there is nearly 10%-20% improvement on the baseline models. This indicates that the CCoE framework can easily boost the base model performance in different downstream domains. To answer the RQ2, Law and medical domains need to have more fine-tuning layers to delivery better test results. We think this is due to that they have less portion in pre-training corpus and some of their knowledge are contemporary. To revise the contemporary knowledge requires to correct the knowledge neurons. Having more fine-tuning layers allows us to modify the knowledge neurons in the lower layers.
@@
The Text-to-SQL requires the LLM to understand the SQL syntax, database schema, and the capability of LLM is also very essential. We observed that there is a performance gap between the models on Spider Leader-board. Comparing to the base model, there is a big improvement. We only conducted limited study on the effect of prompting in Text-to-SQL. In the future, we plan to continue to study the In-Context Learning and calibration of hints to improve the Text-to-SQL performance.
Efficiency Estimation. Table 1 shows the total number of parameters in CCoE 7B-base and 14B-base models. We can observe that the number of activated parameters maintain constantly regardless of the number of experts. For each prompt query, CCoE only needs to activate one of experts to generate the answer, which has the same memory consumption as the base model. The activation ratio can continue to decrease when we scale up the number of experts (Disscussed in Section 4.1). We also can infer the efficiency of training expert. The common approach of enhancing a LLM’s capabilities in 5 domains while preserving its general capabilities requires a mixture training data strategy or multiple stages of training to prevent the forgetting. To estimate the training time for a 7B and 14B model on those 5 domains, we adopt the multiple stages (6 stages) training for 5 domain domains [17]. For the first 5 stages, we train the model on each domain. At the last stage, we mix the data from the 5 domains and the general domain [41]. We understand that this cannot guarantee the model can achieve the same performance as our CCoE models after SFT, but it can provide an approximately estimation on the training procedure of enhancing LLM’s downstream domain capabilities. The CCoE framework uses roughly 25% less training time than training the base models in 5 domains. We think the increase of the training tokens from general domain and the additional stage of training to prevent the catastrophic forgetting are the major factors of increasing training time. Besides, although we conducted multiple stages training for preventing catastrophic forgetting, we still observed there is a big performance degradation on the general abilities and low performance on 5 downstream domains.
We justified the feasibility and efficiency of expert training in CCoE at Section 5.2. To evaluate the current inference efficiency, we calculated the change of TTFT between the CCoE model and base model. We sampled 100 questions with various question length from each domain and evaluate at "round-robin" way. This enable us to evaluate the effectiveness of current expert routing and the inference among collaboration of multiple experts. The evaluation presents that there is nearly no delay during the CCoE inference. Due to the limitation of resources, we fail to conduct more experiments on large group of experts. However, we believe that the delay could slightly increase as we increasing the number of experts. Our evaluations shows that CCoE greatly simplifies the training process and effectively enhancing downstream domain capabilities without hurting each other, and CCoE introduces negligible overheads during expert inference.
@@
Comparison Among Model Families. To answer the RQ 3, We select the open-sourced LLaMA2 7B chat model [42] as our comparison model. LLaMA2 7B has the same model scale as the QWen1.5 7B model. We train the LLaMA2 7B chat model with same number of expert layers and same dataset. From table 4, we can see that LLaMA model is able to receive a higher performance boost than Qwen1.5 base model, but low performance results in 5 domains. We speculate that a comparatively low performance base model can easily to be enhanced through the downstream fine-tuning and data augmentation strategies. Expert training in CCoE requires to utilize the lower layers of the base model to extract the language or even semantic features, which are still decided by the base model’s capabilities. We can conclude that the CCoE framework can boost the performance of base model in different downstream domains regardless of model’s family but expert’s performance is constraint on its base model’s capability. We also realize the scaling law applies to the expert training. It is difficult to continue to improve expert’s performance by increasing the number of training data without changing the expert size.
@@ 6 Summary and Future Work
In this paper, we propose the CCoE framework that aims to boost the base model’s overall performance in all perspectives through utilizing the idea of collaborating multiple of experts (CoE) LLM layers. All of the expert LLM layers are coupled at the base model and becomes the expert sub-network. We conducted the experiments on Math, Code, Law, Medical and Text-to-SQL domains. Our results indicate that there are at nearly 10%-20% improvements on these 5 downstream domains with less training time and resources. This proves that CCoE framework is able to boost the overall performance of the base model with the advantages of low-cost, high-interpretability and efficiency. Moreover, CCoE also has the high extensibility through scaling up the number of expert layers to fulfill more open-ended tasks. However, the current routing of the expert layers depends on the API endpoints trace. In the future, We plan to use the LLM to help us route the prompt to the correct expert sub-network. We can extend the prompt’s content to ask the base model to classify the prompt’s task or domain during the first generation and then run the second inference to route the real prompt to the correct expert sub-network. We understand this could result in the high Time to First Token (TTFT), but there is still great potential of the CCoE network for us to explore.